How would you write the slope-intercept form 2y+3=-1/3(x-2) into standard form? Thank you for your time.
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16946)

2y+3=-1/3(x-2)
multiply by 3
3(2y+3)=x-2
6y+9=x-2
-x+6y+9+2=0
-x+6y+11=0
...
m.ananth@hotmail.ca

Answer by Alan3354(69443)

How would you write the slope-intercept form 2y+3=-1/3(x-2) into standard form?
-------------
That's not slope-int form.
2y+3=-1/3(x-2) = -x/3 + 2/3
2y = -x/3 - 7/3
y = (-1/6)x - 7/6 is slope-int form. Solve for y to get that.
-----------------
y = (-1/6)x - 7/6
6y = -x - 7
x + 6y = -7 in standard form